OpenStudy (anonymous):

Simplify. 7x + 3(x - 2) - 4x + 8 A) 6x + 2 B) 6x + 14 C) 14x + 2 D) 14x + 14 @SkyBudgie9303

OpenStudy (anonymous):

combine like terms will be the first approach

OpenStudy (anonymous):

so what is the oeder of operations? pamdas right?

OpenStudy (anonymous):

order*

OpenStudy (anonymous):

pemdas*

OpenStudy (anonymous):

Perenthises Exponets Multiply Divide Add Subtract

OpenStudy (anonymous):

anyways... i disagree. i think you should do the parenthesis first

OpenStudy (anonymous):

no need for the attitude

OpenStudy (anonymous):

first you must solve 3(x-2) and then combine like terms

OpenStudy (anonymous):

Ok so 3(x) =3x

OpenStudy (anonymous):

Is that right?

OpenStudy (anonymous):

https://www.youtube.com/watch?v=zanq7gmXY88 this is Flocabulary order of operations

OpenStudy (anonymous):

im homeschooled

OpenStudy (anonymous):

me 2 @THEHELPER123

OpenStudy (anonymous):

?!?!? what school

OpenStudy (anonymous):

keystone high school :)

OpenStudy (anonymous):

= 7x + 3(x - 2) - 4x + 8 = 7x + 3x - 6 - 4x + 8 = 7x - 4x + 3x + 8 - 6 = 6x + 2

OpenStudy (anonymous):

im in middle school :(

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Oh ok let me write this down

OpenStudy (anonymous):

whats ur grade lvl @nono266

OpenStudy (anonymous):

10th

OpenStudy (anonymous):

2 digits

OpenStudy (anonymous):

im in 6th

OpenStudy (anonymous):

Ok got it. Thank you!